OpenStudy (anonymous):

*MEDAL*FAN* Create a system of equations that includes one linear equation and one quadratic equation. If you can, please help me to know how to solve it algebraically and how to graph it. Thanks! :D @Lady.Liv1776 @bibby @paki @Preeta @TheSmartOne

OpenStudy (anonymous):

@sammixboo Please help me

OpenStudy (anonymous):

hmmmmm

OpenStudy (anonymous):

x + 1)(x – 3) = 0 x + 1 = 0 or x – 3 = 0 x = –1 or x = 3

OpenStudy (anonymous):

and

OpenStudy (anonymous):

Is that linear? or quadratic

OpenStudy (anonymous):

quad

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

4 X - 32 = 8

OpenStudy (anonymous):

linear

OpenStudy (anonymous):

your welcome